At our fishing club, each person records the weight of the largest fish caught for the day and the mean weight is calculated so that members can see how they are going compared to other fishermen. The other day. the whole club except for one person was below the average. How is this possible?
Answer by fractalier(6550)

The one person must have caught a monster.
Let's say he caught a 200-lb fish.
If four others caught ten-pounders, the mean would be
(200+10+10+10+10)/5 = 240/5 = 48 lbs.